# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ 3.2.3 Content retrieval and recommendation services
Generative retrieval (GR) was pioneered by Metzler et al. [90]. Unlike traditional retrieval, which adheres to the “index-retrieve-then-rank” paradigm, the GR paradigm employs a single model to obtain results from query input. The model generally involves deep-learning based transformers, generating output token-by-token. More recent work in [95] introduces learning-to-rank training to enhance the performance system up to 30%. GR has several advantages including substituting the bulky external index with an internal index (i.e., model parameters), significantly reducing memory usage, and enabling optimization during end-to-end model training towards a universal objective for information retrieval tasks. Conversational question answering techniques have been integrated to enhance the document retrieval [94].
When retrieving visual content, recent work exploits generative models to enhance content-based model search [92]. These models decode the text, image, or video query into samples of possible outputs, which are then used to learn statistics for better matching between the query and output candidates. DMs are also employed for visual retrieval tasks, where they learn joint data distributions between text queries and video candidates [96]. A comprehensive survey on Generative Information Retrieval is available in [272].
While the retrieval task involves users directly defining a specific query input, recommendation services operate by retrieving content based on previous usage patterns. Essentially, a recommendation engine is a system that suggests products, services, or information to users through data analysis. Research in [273] has reported a positive association between buyers’ attitudes toward AI and their behavioral intention to accept AI-based recommendations, with potential for further growth. Notable examples include the recommendation framework developed by Google [93], which utilizes GR. This framework assigns Semantic IDs to each item and trains a retrieval model to predict the Semantic ID of an item that a given user may engage with. A report by Aggarwal et al. [274] states that the recommendation accuracy of recommendation services has increased from 45.0% to 91.5% with the integration of generative AI.
@@ 3.2.4 Intelligent assistants
Current LLMs obviously enhance the performance of intelligent assistants, designed to understand complex inquiries and generate more natural conversational responses, such as Sasha [97]. Generative AI can also be used to enhance the performance of human customer support agents, aiding in search and summarization, as discussed in the previous section. Brynjolfsson et al. [275] examined the implementation of a generative AI tool designed to offer conversational guidance to customer support agents. Their research revealed that AI assistance significantly enhances problem resolution and customer satisfaction. Furthermore, they observed that AI recommendations prompt low-skill workers to adopt communication styles akin to those of high-skill workers. AI-based intelligent assistants may currently be more focused on educational purposes, but they can clearly help artists write more efficiently [276] or assist in customizing personal requirements [277]. The performance of personalized assistants can be enhanced with domain-specific knowledge to provide more in-depth responses to users [278].
@@ 3.3.1 Enhancement
In our previous review paper [8], we discussed AI technologies for contrast enhancement and colorization as separate topics, as methods were developed specifically for each task. However, in recent years, there has been a shift towards addressing more complex issues, such as those encountered in low-light environments and underwater scenarios. These real-world situations often involve a combination of challenges, including low contrast, color imbalance, and noise.
@@
In low-light conditions, scenes often exhibit low contrast, leading to focusing difficulties or the need for long exposures, which can result in blurred images and videos. To address this, LEDNet [279] has introduced a synthetic dataset for such scenarios and incorporated a learnable non-linear activation function within the network to enhance feature intensities. Meanwhile, SNR-Aware [98] estimates spatial-varying Signal-to-Noise Ratio (SNR) maps and proposes local and global learning branches using ResNet and transformer architectures, respectively. NeRCo [108] addresses the low-light problem with INR, which unifies the diverse degradation factors of real-world scenes with a controllable fitting function. Diffusion models (DMs) have also become popular choices for low-light image enhancement [104, 105, 106]. Diff-Retinex [105] formulates the low-light image enhancement problem into Retinex decomposition, and employs multi-path generative diffusion networks to reconstruct the normal-light Retinex probability distribution. A recent state-of-the-art approach presented in [106] decomposes images into high and low frequencies using wavelet transform. High frequencies are enhanced using a transformer-based pipeline, while the low frequencies undergo a diffusion process. This method achieves nearly 2.8dB improvement over the state-of-the-art transformer-based approach, e.g. LLFormer [100], and significantly better than INR-based method, NeRCo [108], on a real low-light image benchmarking dataset. The technique has been extended for video enhancement in [107]. The output of the enhancement typically depends on user preferences. This has been viewed as a one-to-many inverse problem, with attempts to solve it using Bayesian approaches. For example, a Bayesian Enhancement Model (BEM) [280] incorporates Bayesian Neural Networks (BNNs) to capture data uncertainty and produce diverse outputs. The method can be used with Transformers or Mamba as the architecture backbone.
@@
Regarding video enhancement, transformer and DMs are still in their early stages. STA-SUNet [101] has demonstrated that using transformers for low-light video enhancement outperforms CNN-based methods [281]. The recent Mamba-based network [282] also demonstrates promising results, outperforming STA-SUNet by more than 2 dB in PSNR. It is important to note that low-light enhancement is subjective. While most training datasets use normal lighting conditions as ground truth [283], the enhanced images and videos may alter the mood and tone of the content. Therefore, the tools for creative industries should be adjustable, not only for entire images and videos but also adaptive to specific areas and content. For instance, CLE Diffusion [129] enables user-friendly editing of lighting with fine-grained regional controllability.
@@
Recent efforts have focused on enhancing User-Generated Content (UGC) videos—authentic recordings created by individuals rather than brands, often showcasing real experiences with products or services. The winning solution of the NTIRE 2025 Challenge on UGC Video Enhancement [284] implemented a pipeline of four sequential modules: color enhancement, denoising, BasicVSR++ restoration [285], and SwinIR [115]. This method achieved a 17% higher subjective score than the second-place entry, which used a two-stage framework, highlighting a notable improvement in perceived visual quality.
@@ 3.3.2 Style transfer
StyTr2 [109] is the first transformer-based method for style transfer, applying content as a query and style as a key of attention. InST [112] utilizes Stable Diffusion Models as the generative backbone and introduces an attention-based textual inversion module to learn the description of the content. StableVideo [113] uses a text prompt to describe the desired appearance of the output, transforming the input video to have a new look based on a diffusion model. For instance, a video of a white car driving in summer can be altered to show a red car driving in winter. A large pre-trained DM is employed in [111], where the style is injected to manipulate the self-attention of the decoder. To deal with the disharmonious color, they propose an adaptive instance normalization. A survey of style transfer using transformers and diffusion models can be found in [286]. Implicit Neural Representations (INRs) are less commonly used in style transfer tasks due to the difficulty of modeling the cross-representation between style and content. Moon et al. [110] combined INRs with vision transformers for generalizable style transfer; however, the results remain limited in quality. In contrast, the method proposed by Kim et al. [114] uses multilayer perceptrons (MLPs) to map image coordinates to the colors of the stylized output, guided by features extracted from both the content and style inputs to allow controllability.
@@ 3.3.3 Upscaling imagery: super-resolution (SR)
The first use of a transformer, called ESRT, was for capturing long-term dependencies, such as repeating patterns in buildings. This was done in the feature domain extracted by a lightweight CNN module [116], outperforming those that use only CNNs. Since then, most SISR methods have been based on transformers. The Hybrid Attention Transformer (HAT) [118] was introduced, which improves the SR quality over ESRT by more than 2dB when upscaling 2$\times$-4$\times$. However, the NTIRE 2023 Real-Time Super-Resolution Challenge [287] showed that the winner, Bicubic++ [288], uses only convolutional layers and achieves the fastest speed at 1.17ms in upscaling 720p to 4K images. This method is significantly faster than any of the participants in the NTIRE 2025 Challenge [289], where Transformer-based architectures continue to dominate as the mainstream approach.
For DMs, SR3 by Google [123] has produced truly impressive results. It operates by learning to transform a standard normal distribution into an empirical data distribution through a sequence of refinement steps, interpolating in a cascaded manner—upscaling 4$\times$ at a time. Later, IDM [125] combines INR with a U-Net denoising model in the reverse process of the DM. It is crucial to emphasize again that DMs are generative models. The SR results are generated based on the statistics we provide to the model during training (LR training samples). This is not for a restoration task, but rather for synthetic generation. A survey in SISR using DMs can be found in [290].
For video SR, numerous methods have emerged as part of a unified enhancement framework, as discussed in the previous section. One of the pioneering works to incorporate transformers specifically for video SR tasks is the Trajectory-aware Transformer for Video Super-Resolution (TTVSR) [117]. Although the results are slightly inferior to those of BasicVSR++ [285], which employs CNN and was introduced around the same time, both methods significantly enhance detail and sharpness compared to previous approaches, albeit not in real time. To address this limitation, the Deformable Attention Pyramid [291] has been introduced, offering slightly lower quality but a speed-up of over 3$\times$. Recently, Adobe announced their VideoGigaGAN [121], which can perform 8$\times$ upsampling. This is achieved by adding flow estimation and temporal self-attention to the GigaGAN upsampler [120], which is primarily used for image SR, and text-to-image synthesis. Cao et al. [126] introduce a zero-shot video super-resolution framework that leverages a pre-trained image diffusion model, and replaces the spatial self-attention layer with a novel short-long-range (SLR) temporal attention layer. Recently, SeedVR integrated text information (captions) into a Diffusion Transformer (DiT) model, achieving state-of-the-art performance in video super-resolution.
@@ 3.3.4 Restoration
In our previous review paper [8], we categorized the work on restoration into several different types of distortions, including deblurring, denoising, dehazing, and mitigating atmospheric turbulence. Recent work, however, uses a unified network architecture to address these as inverse problems $y=hx+n$, where $x$ and $y$ are the ideal and observed data, respectively. $h$ is a degradation function, such as blur, and $n$ is additive noise. Often, the super-resolution task is also considered as an inverse problem, meaning $h$ includes the downsampling process. Note that although designed as a single network, the model is trained with each distorted dataset separately.
@@
The pioneering transformer-based method for image restoration, SwinIR [115], employs several concatenated Swin Transformer blocks [16]. SwinIR surpasses state-of-the-art CNN-based methods proposed up to the year 2021 in super-resolution and denoising tasks. The model is smaller and reconstructs fine details more effectively. Other two popular approaches that emerged in the same timeframe are Uformer [130] and Restormer [131]. Both incorporate Transformer blocks into hierarchical encoder-decoder networks, employing skip connections similar to those in U-Net. Their objective was to restore noisy images, sharpen blurry images, and remove rain. The networks focused on predicting the residual $R$ and obtaining the restored image $\hat{x}$ through $\hat{x}=y+R$. While their performance is very similar, Restormer has half the parameters of Uformer. More recently, GRL by Li et al. [119] exploits a hierarchy of features in a global, regional, and local range using different ways to compute self-attentions as an image often show similarity within itself in different scales and areas, which outperforms SwinIR and Restormer. Additionally, Fei et al. introduced the Generative Diffusion Prior [128] for unsupervised learning, aiming to model posterior distributions for image restoration and enhancement. VmambaIR [143] incorporates Mamba blocks into the U-Net architecture, achieving superior performance compared to SwinIR and Restormer in both visual quality and model size.
@@
For video restoration, the general framework comprises frame alignment, feature fusion and reconstruction. The process could be similar to image restoration but input multiple frames and run through the sequences in a sliding window manner to exploit the temporal information within a number of consecutive frames. Recently, Video Restoration Transformer (VRT) [103] and its improved version with a recurrent process (RVRT) [99], have emerged as the state of the arts for video super-resolution, deblurring, denoising, and frame interpolation. This method introduces temporal reciprocal self-attention in the transformer architecture and parallel warping using MLP. These innovations enable parallel computation and outperform the previous state-of-the-art methods by up to 2.16dB on benchmark datasets. FMA-Net [102] proposed multi-attention for joint video super-resolution and deblurring, achieving fast runtime with nearly 40% improvement over RVRT, and the restored quality was reported better by up to 3%.
@@
For audio restoration, most software discussed in Section 3.1.2 offers tools for enhancing audio quality, such as eliminating background noise, echo, microphone rumble, and occasionally room reverberation, which have been well-established even before the advent of deep learning. There have been efforts to utilize AI for learning global contextual information to aid in the removal of unwanted sounds, leading to better final quality [134]. The latest advancements in this domain are primarily focused on addressing issues where significant portions of the audio data are missing. For instance, Moliner et al. [124] tackle problems such as audio bandwidth extension, inpainting, and declipping by treating them as inverse problems using a diffusion model. For a comprehensive survey on the use of diffusion models in restoration tasks, refer to [292].
@@
i) Deblurring: A lightweight deep CNN model was recently proposed in [293], where a new discriminative temporal feature fusion has been introduced to select the most useful spatial and temporal features from adjacent frames. Feature propagation along the video is done in the wavelet domain. The deblurring performance is comparable to RVRT [99], but it is 5 times faster. DaBiT [133] mitigates focal blur content with depth information and applies SR for further enhancing fine details. Note that not only in software, but AI technologies have also been integrated into hardware. This includes autofocus, which is crucial for capturing sharp images of subjects, especially in dynamic environments where manual adjustments are impractical due to rapid movement. AI-driven autofocus methods have emerged, often tailored for specific camera hardware. For instance, Choi et al. proposed an autofocus model optimized for dual-pixel Canon cameras [294]. Additionally, Yang et al. investigated the correlation between language input and blur map estimation, utilizing semantic cues to enhance autofocus performance [132]. Remarkably, their model achieves comparable results to previous state-of-the-art methods while being more lightweight [295]. Autofocus could be used in conjunction with real-time object tracking (see Section 3.4.3) to produce desirable sharpness for moving objects in the video. Recently, Feng et al. [147] proposed a novel residual diffusion deblurring framework that integrates a conditional diffusion model guided by a defocus map and incorporates residual learning into the single-image defocus deblurring process.
@@
ii) Denoising: SUNet [18] applies Swin transformer blocks combined in a UNet-like architecture. Denoising with diffusion models (DMs) [144] has been proposed by diffusing with estimated noise that is closer to real-world noise rather than Gaussian noise, achieving better performance than SwinIR [115] and Uformer [130]. INR with complex Gabor wavelets as activation functions show promising denoising results [42]. The NTIRE 2025 Image Denoising Challenge [296] revealed that the top-performing methods combined transformer-based and convolutional network architectures. Similarly, recent advances in video denoising also adopt a hybrid approach that integrates both architectures [142, 141].
@@
iii) Dehazing: Vision transformers for single image dehazing were proposed in DehazeFormer [135]. Similar to SUNet, it is a UNet-like architecture, but introduces Rescale Layer Normalization for better suit on improving contrast. The Fast Fourier Transform (FFT) has been employed in [140] due to the phase spectrum conveying more structural detail than the amplitude spectrum and demonstrating greater robustness to contrast distortion and noise. Then cross-attention between the RGB and YCbCr color spaces is applied. This approach achieves nearly 5 dB higher PSNR than DehazeFormer on a real-world smoke dataset. For video dehazing, Xu et al. [136] introduced a recurrent multi-range scene radiance recovery module with the space-time deformable attention. They also employ physics prior to inform haze attenuation. This method outperforms DehazeFormer by approximately 1dB.
@@
iv) Mitigating atmospheric turbulence: Similar to dehazing, physics-inspired models have been widely developed to remove turbulence distortion [146, 148], while complex-valued CNNs have been proposed to exploit phase information [297]. There was also an attempt to use instance normalization (INR) to solve this problem, offering tile and blur correction [148]. However, diffusion models outperform on a single image [145], and transformer-based methods remain state-of-the-art for restoring videos [138, 139]. Mamba architecture, employed in [Hill2025MAMAT], outperforms Transformers and improves object detection performance. A recent review can be found in [299].

### v5

@@ 3.2.3 Content retrieval and recommendation services
Generative retrieval (GR) was pioneered by Metzler et al. [Metzler:Rethinking:2021]. Unlike traditional retrieval, which adheres to the “index-retrieve-then-rank” paradigm, the GR paradigm employs a single model to obtain results from query input. The model generally involves deep-learning based transformers, generating output token-by-token. More recent work in [li2024learning] introduces learning-to-rank training to enhance the performance system up to 30%. GR has several advantages including substituting the bulky external index with an internal index (i.e., model parameters), significantly reducing memory usage, and enabling optimization during end-to-end model training towards a universal objective for information retrieval tasks. Conversational question answering techniques have been integrated to enhance the document retrieval [li:unigen:2024].
When retrieving visual content, recent work exploits generative models to enhance content-based model search [Lu:content:2023]. These models decode the text, image, or video query into samples of possible outputs, which are then used to learn statistics for better matching between the query and output candidates. DMs are also employed for visual retrieval tasks, where they learn joint data distributions between text queries and video candidates [Jin:DiffusionRet:2023]. A comprehensive survey on Generative Information Retrieval is available in [Li:From:2025].
While the retrieval task involves users directly defining a specific query input, recommendation services operate by retrieving content based on previous usage patterns. Essentially, a recommendation engine is a system that suggests products, services, or information to users through data analysis. Research in [CHUA:AI:2023] has reported a positive association between buyers’ attitudes toward AI and their behavioral intention to accept AI-based recommendations, with potential for further growth. Notable examples include the recommendation framework developed by Google [Rajput:recommender:2023], which utilizes GR. This framework assigns Semantic IDs to each item and trains a retrieval model to predict the Semantic ID of an item that a given user may engage with. A report by Aggarwal et al. [aggarwal2025evolution] states that the recommendation accuracy of recommendation services has increased from 45.0% to 91.5% with the integration of generative AI.
@@ 3.2.4 Intelligent assistants
Current LLMs obviously enhance the performance of intelligent assistants, designed to understand complex inquiries and generate more natural conversational responses, such as Sasha [King:Sasha:2024]. Generative AI can also be used to enhance the performance of human customer support agents, aiding in search and summarization, as discussed in the previous section. Brynjolfsson et al. [brynjolfsson:generative:2023] examined the implementation of a generative AI tool designed to offer conversational guidance to customer support agents. Their research revealed that AI assistance significantly enhances problem resolution and customer satisfaction. Furthermore, they observed that AI recommendations prompt low-skill workers to adopt communication styles akin to those of high-skill workers. AI-based intelligent assistants may currently be more focused on educational purposes, but they can clearly help artists write more efficiently [Lee:design:2024] or assist in customizing personal requirements [sajja2024ai]. The performance of personalized assistants can be enhanced with domain-specific knowledge to provide more in-depth responses to users [jiang2025domain].
@@ 3.3.1 Enhancement
In our previous review paper [Anantrasirichai:AI:2022], we discussed AI technologies for contrast enhancement and colorization as separate topics, as methods were developed specifically for each task. However, in recent years, there has been a shift towards addressing more complex issues, such as those encountered in low-light environments and underwater scenarios. These real-world situations often involve a combination of challenges, including low contrast, color imbalance, and noise.
@@
In low-light conditions, scenes often exhibit low contrast, leading to focusing difficulties or the need for long exposures, which can result in blurred images and videos. To address this, LEDNet [Zhou:LEDNet:2022] has introduced a synthetic dataset for such scenarios and incorporated a learnable non-linear activation function within the network to enhance feature intensities. Meanwhile, SNR-Aware [Xu:SNR:2022] estimates spatial-varying Signal-to-Noise Ratio (SNR) maps and proposes local and global learning branches using ResNet and transformer architectures, respectively. NeRCo [Yang:Implicit:2023] addresses the low-light problem with INR, which unifies the diverse degradation factors of real-world scenes with a controllable fitting function. Diffusion models (DMs) have also become popular choices for low-light image enhancement [HOU:Global:2023, Yi:Diff:2023, Jiang:Low:2023]. Diff-Retinex [Yi:Diff:2023] formulates the low-light image enhancement problem into Retinex decomposition, and employs multi-path generative diffusion networks to reconstruct the normal-light Retinex probability distribution. A recent state-of-the-art approach presented in [Jiang:Low:2023] decomposes images into high and low frequencies using wavelet transform. High frequencies are enhanced using a transformer-based pipeline, while the low frequencies undergo a diffusion process. This method achieves nearly 2.8dB improvement over the state-of-the-art transformer-based approach, e.g. LLFormer [Wang:Ultra:2023], and significantly better than INR-based method, NeRCo [Yang:Implicit:2023], on a real low-light image benchmarking dataset. The technique has been extended for video enhancement in [lin2024lowlight]. The output of the enhancement typically depends on user preferences. This has been viewed as a one-to-many inverse problem, with attempts to solve it using Bayesian approaches. For example, a Bayesian Enhancement Model (BEM) [huang2025bayesian] incorporates Bayesian Neural Networks (BNNs) to capture data uncertainty and produce diverse outputs. The method can be used with Transformers or Mamba as the architecture backbone.
@@
Regarding video enhancement, transformer and DMs are still in their early stages. STA-SUNet [Lin:SPATIO:2024] has demonstrated that using transformers for low-light video enhancement outperforms CNN-based methods [anantrasirichai:BVI:2024]. The recent Mamba-based network [huang2025bvi] also demonstrates promising results, outperforming STA-SUNet by more than 2 dB in PSNR. It is important to note that low-light enhancement is subjective. While most training datasets use normal lighting conditions as ground truth [Lin:BVI-RLV:2024], the enhanced images and videos may alter the mood and tone of the content. Therefore, the tools for creative industries should be adjustable, not only for entire images and videos but also adaptive to specific areas and content. For instance, CLE Diffusion [Yin:CLE:2023] enables user-friendly editing of lighting with fine-grained regional controllability.
@@
Recent efforts have focused on enhancing User-Generated Content (UGC) videos—authentic recordings created by individuals rather than brands, often showcasing real experiences with products or services. The winning solution of the NTIRE 2025 Challenge on UGC Video Enhancement [safonov2025ntire] implemented a pipeline of four sequential modules: color enhancement, denoising, BasicVSR++ restoration [Chan:BasicVSR:2022], and SwinIR [Liang:SwinIR:2021]. This method achieved a 17% higher subjective score than the second-place entry, which used a two-stage framework, highlighting a notable improvement in perceived visual quality.
@@ 3.3.2 Style transfer
StyTr2 [Deng:StyTr2:2022] is the first transformer-based method for style transfer, applying content as a query and style as a key of attention. InST [Zhang:Inversion:2023] utilizes Stable Diffusion Models as the generative backbone and introduces an attention-based textual inversion module to learn the description of the content. StableVideo [Chai:StableVideo:2023] uses a text prompt to describe the desired appearance of the output, transforming the input video to have a new look based on a diffusion model. For instance, a video of a white car driving in summer can be altered to show a red car driving in winter. A large pre-trained DM is employed in [Chung_2024_CVPR], where the style is injected to manipulate the self-attention of the decoder. To deal with the disharmonious color, they propose an adaptive instance normalization. A survey of style transfer using transformers and diffusion models can be found in [ZHOU:Bridging:2025]. Implicit Neural Representations (INRs) are less commonly used in style transfer tasks due to the difficulty of modeling the cross-representation between style and content. Moon et al. [Moon:generalizable:2023] combined INRs with vision transformers for generalizable style transfer; however, the results remain limited in quality. In contrast, the method proposed by Kim et al. [Kim:Controllable:2024] uses multilayer perceptrons (MLPs) to map image coordinates to the colors of the stylized output, guided by features extracted from both the content and style inputs to allow controllability.
@@ 3.3.3 Upscaling imagery: super-resolution (SR)
[…]
@@ 3.3.4 Restoration
In our previous review paper [Anantrasirichai:AI:2022], we categorized the work on restoration into several different types of distortions, including deblurring, denoising, dehazing, and mitigating atmospheric turbulence. Recent work, however, uses a unified network architecture to address these as inverse problems $y=hx+n$, where $x$ and $y$ are the ideal and observed data, respectively. $h$ is a degradation function, such as blur, and $n$ is additive noise. Often, the super-resolution task is also considered as an inverse problem, meaning $h$ includes the downsampling process. Note that although designed as a single network, the model is trained with each distorted dataset separately.
@@
The pioneering transformer-based method for image restoration, SwinIR [Liang:SwinIR:2021], employs several concatenated Swin Transformer blocks [Liu:Swin:2021]. SwinIR surpasses state-of-the-art CNN-based methods proposed up to the year 2021 in super-resolution and denoising tasks. The model is smaller and reconstructs fine details more effectively. Other two popular approaches that emerged in the same timeframe are Uformer [Wang:Uformer:2022] and Restormer [Zamir:Restormer:2022]. Both incorporate Transformer blocks into hierarchical encoder-decoder networks, employing skip connections similar to those in U-Net. Their objective was to restore noisy images, sharpen blurry images, and remove rain. The networks focused on predicting the residual $R$ and obtaining the restored image $\hat{x}$ through $\hat{x}=y+R$. While their performance is very similar, Restormer has half the parameters of Uformer. More recently, GRL by Li et al. [li:GRL:2023] exploits a hierarchy of features in a global, regional, and local range using different ways to compute self-attentions as an image often show similarity within itself in different scales and areas, which outperforms SwinIR and Restormer. Additionally, Fei et al. introduced the Generative Diffusion Prior [Fei:Generative:2023] for unsupervised learning, aiming to model posterior distributions for image restoration and enhancement. VmambaIR [Shi:VmambaIR:2025] incorporates Mamba blocks into the U-Net architecture, achieving superior performance compared to SwinIR and Restormer in both visual quality and model size.
@@
For video restoration, the general framework comprises frame alignment, feature fusion and reconstruction. The process could be similar to image restoration but input multiple frames and run through the sequences in a sliding window manner to exploit the temporal information within a number of consecutive frames. Recently, Video Restoration Transformer (VRT) [Liang:VRT:2024] and its improved version with a recurrent process (RVRT) [liang:RVRT:2022], have emerged as the state of the arts for video super-resolution, deblurring, denoising, and frame interpolation. This method introduces temporal reciprocal self-attention in the transformer architecture and parallel warping using MLP. These innovations enable parallel computation and outperform the previous state-of-the-art methods by up to 2.16dB on benchmark datasets. FMA-Net [Youk:FMA:2024] proposed multi-attention for joint video super-resolution and deblurring, achieving fast runtime with nearly 40% improvement over RVRT, and the restored quality was reported better by up to 3%.
@@
For audio restoration, most software discussed in Section 3.1.2 offers tools for enhancing audio quality, such as eliminating background noise, echo, microphone rumble, and occasionally room reverberation, which have been well-established even before the advent of deep learning. There have been efforts to utilize AI for learning global contextual information to aid in the removal of unwanted sounds, leading to better final quality [Yu:DBT:2022]. The latest advancements in this domain are primarily focused on addressing issues where significant portions of the audio data are missing. For instance, Moliner et al. [Moliner:solving:2023] tackle problems such as audio bandwidth extension, inpainting, and declipping by treating them as inverse problems using a diffusion model. For a comprehensive survey on the use of diffusion models in restoration tasks, refer to [10902142].
@@
i) Deblurring: A lightweight deep CNN model was recently proposed in [Pan:Deep:2023], where a new discriminative temporal feature fusion has been introduced to select the most useful spatial and temporal features from adjacent frames. Feature propagation along the video is done in the wavelet domain. The deblurring performance is comparable to RVRT [liang:RVRT:2022], but it is 5 times faster. DaBiT [Morris:DaBiT:2024] mitigates focal blur content with depth information and applies SR for further enhancing fine details. Note that not only in software, but AI technologies have also been integrated into hardware. This includes autofocus, which is crucial for capturing sharp images of subjects, especially in dynamic environments where manual adjustments are impractical due to rapid movement. AI-driven autofocus methods have emerged, often tailored for specific camera hardware. For instance, Choi et al. proposed an autofocus model optimized for dual-pixel Canon cameras [Choi:Exploring:2023]. Additionally, Yang et al. investigated the correlation between language input and blur map estimation, utilizing semantic cues to enhance autofocus performance [yang:ldp:2023]. Remarkably, their model achieves comparable results to previous state-of-the-art methods while being more lightweight [Yang:K3DN:2023]. Autofocus could be used in conjunction with real-time object tracking (see Section 3.4.3) to produce desirable sharpness for moving objects in the video. Recently, Feng et al. [feng2025residual] proposed a novel residual diffusion deblurring framework that integrates a conditional diffusion model guided by a defocus map and incorporates residual learning into the single-image defocus deblurring process.
@@
ii) Denoising: SUNet [Fan:SUNet:2022] applies Swin transformer blocks combined in a UNet-like architecture. Denoising with diffusion models (DMs) [yang:realworld:2023] has been proposed by diffusing with estimated noise that is closer to real-world noise rather than Gaussian noise, achieving better performance than SwinIR [Liang:SwinIR:2021] and Uformer [Wang:Uformer:2022]. INR with complex Gabor wavelets as activation functions show promising denoising results [Saragadam:wire:2023]. The NTIRE 2025 Image Denoising Challenge [sun2025tenth] revealed that the top-performing methods combined transformer-based and convolutional network architectures. Similarly, recent advances in video denoising also adopt a hybrid approach that integrates both architectures [Jin:Masked:2025, Yue:RViDeformer:2025].
@@
iii) Dehazing: Vision transformers for single image dehazing were proposed in DehazeFormer [Song:vision:2023]. Similar to SUNet, it is a UNet-like architecture, but introduces Rescale Layer Normalization for better suit on improving contrast. The Fast Fourier Transform (FFT) has been employed in [fang2025guided] due to the phase spectrum conveying more structural detail than the amplitude spectrum and demonstrating greater robustness to contrast distortion and noise. Then cross-attention between the RGB and YCbCr color spaces is applied. This approach achieves nearly 5 dB higher PSNR than DehazeFormer on a real-world smoke dataset. For video dehazing, Xu et al. [Xu:Video:2023] introduced a recurrent multi-range scene radiance recovery module with the space-time deformable attention. They also employ physics prior to inform haze attenuation. This method outperforms DehazeFormer by approximately 1dB.
@@
iv) Mitigating atmospheric turbulence: Similar to dehazing, physics-inspired models have been widely developed to remove turbulence distortion [Jaiswal:Physics:2023, Jiang:NeRT:2023], while complex-valued CNNs have been proposed to exploit phase information [Atmospheric:2023]. There was also an attempt to use instance normalization (INR) to solve this problem, offering tile and blur correction [Jiang:NeRT:2023]. However, diffusion models outperform on a single image [Nair:AT-DDPM:2023], and transformer-based methods remain state-of-the-art for restoring videos [Zhang:Image:2024, zou2024deturb]. Mamba architecture, employed in [Hill2025MAMAT], outperforms Transformers and improves object detection performance. A recent review can be found in [Hill2025].